OpenStudy (anonymous):

If cos(x) = 2/3, then the value of cos(2x) is a) -1/9 b) -5/9 c) 5/9 d) 1/9 Is it (a) or (d)?

OpenStudy (anonymous):

\[2\cos^2(x)-1\]

OpenStudy (anonymous):

so \[2(\frac{2}{3})^2-1=2\times \frac{4}{9}-1=\frac{8}{9}-1=-\frac{1}{9}\]

OpenStudy (anonymous):

Great! Thanks! That's what I was thinking, I just subtracted wrong.